OpenStudy (anonymous):

In a 4 degree -45 degree -90 degree triangle, the length of the hypotenuse is 11. Find the length of one of the legs ? I need help please is explain how to do this

OpenStudy (zephyr141):

so basically that?

OpenStudy (zephyr141):

if that's the triangle you're thinking of then here's how to approach it. first remember to use the Pythagorean theorem.\[a^2+b^2=c^2\]where a and b are the sides (base and height) and c is the hypotenuse. then you have to remember your trig functions. like SOH CAH TOA and stuff. we have the angles of the triangle so we can do this. lets start with the bottom length (base). so we have the angle 45 degree there. now to find the base you use cosine. remember cosine is adjacent over hypotenuse. so we can make an equation using that. \[\cos(45)=\frac{ adjacent }{ 11 }\]and to find our adjacent length (or base of our triangle) we just work the equation to be equal to the unknown.\[11*\cos(45)=adjacent\]and that's how you find a length of the triangle. well we didn't use the Pythagorean theorem but it's always good to remember so yeah.